### v1

Abstract
Diffusion Language Models (DLMs) offer a promising parallel generation paradigm but suffer from slow inference due to numerous refinement steps and the inability to use standard KV caching. We introduce CDLM (Consistency Diffusion Language Models), a training-based acceleration method that simultaneously tackles both bottlenecks. CDLM integrates consistency modeling to drastically reduce the number of required sampling steps by enabling multi-token finalization. Furthermore, we enforce a block-wise causal attention mask during fine-tuning, making the model fully compatible with KV caching. Experiments show CDLM achieves $3.6\times$$14.5\times$ lower latency while maintaining competitive accuracy on math and coding tasks. The full training and evaluation code is available at https://github.com/SqueezeAILab/CDLM.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Large Language Models (LLMs) have triggered a paradigm shift in AI, excelling at tasks such as instruction following, multi-turn dialogue, and code generation openai2023gpt4. Most current LLMs are autoregressive, decoder-only Transformers trained to predict the next token under a causal attention mask dubey2024llama; zhang2022opt. While training is parallelizable via teacher forcing, inference inherits a rigid sequential dependency, becoming a key efficiency bottleneck. Moreover, such models cannot exploit bidirectional context, which limits performance on tasks such as text infilling, editing, and global planning hu2024bai; song2025seeddiffusion.
As a promising alternative, diffusion language models (DLMs) have drawn attention for their potential in text generation. Inspired by the success of diffusion models in image, audio, and video synthesis ho2020ddpm, DLMs iteratively denoise random or masked token sequences into coherent text gong2023diffuseq; li2022diffusionlm. This process updates all token positions in parallel per step, removing the token-level sequential dependency inherent to autoregressive decoding. Closed-source models, primarily in code generation, report up to $10\times$ higher throughput than autoregressive models while preserving quality deepmind2025geminidiffusion; khanna2025mercury; song2025seeddiffusion. However, open-source DLMs remain significantly slower nie2025large; ye2025dream7b, largely because (1) bidirectional attention prohibits standard KV caching and (2) they demand many refinement steps, often proportional to sequence length, to reach high fidelity kim2025beyond.
Many recent works aim to accelerate DLM inference. One primary strategy leverages block-wise decoding with KV caching, either via approximate caching with periodic updates hu2025freecache; liu2025dllmcache; ma2025dkvcache; wu2025fastdllm or by fine-tuning the model into a fully block-wise causal architecture so that caching becomes natural wang2025d2f; wu2025fastdllmv2. The other major line of work focuses on parallel sampling algorithms, which effectively decode multiple tokens per denoising step and reduce the total number of iterations benhamu2025ebsampler; wu2025fastdllm.
Turning to diffusion-based applications in vision domains, many works have successfully reduced denoising to few-step or even one-step generation yin2024dmd; yin2025causvid. A prominent technique in that realm is consistency modeling: consistency models are trained to minimize discrepancies between adjacent denoising states along the probability flow ODE, enabling near one-step generation song2023consistency; song2023improved. Although originally developed for continuous data, the consistency paradigm has begun to influence discrete settings, including parallel decoding for autoregressive language models kou2024cllms and masked diffusion frameworks xu2025unicms, indicating its strong promise in discrete language domains.
To drive DLMs to converge to coherent text in far fewer refinement steps, we present CDLM, a training-based acceleration method that integrates Consistency modeling into Diffusion Language Models. Concretely, the student is trained to finalize multiple tokens per step by distilling from a fully bidirectional teacher and by aligning its predictions between a less-informed state and its block-completion state along the decoding trajectory. We impose a block-wise causal attention mask and jointly optimize three objectives. At inference time, the model decodes sequentially over blocks while applying confidence-thresholded parallel decoding within each block. This design directly tackles two core bottlenecks of current DLMs: (i) excessive sampling steps (mitigated by distillation and consistency) and (ii) cache inefficiency (removed by block-wise causality). In particular, we make the following contributions:
  • Consistency-guided distillation. We bring consistency modeling to DLMs by defining a token-level decoding trajectory and coupling it with distillation from a fully bidirectional teacher. This joint objective provides explicit supervision for multi-token finalization while keeping predictions stable within each block (see Sections 3 and 4.2).
  • Cache-friendly block causality. By fine-tuning with a block-wise causal attention mask (Figure 2), we enable bidirectional DLMs to support block-wise KV caching and early stopping at block boundaries (see Sections 4.1 and 4.3).
  • End-to-end speedups. With 8 hours of training for Dream and 16 hours for LLaDA, CDLM cuts refinement steps by $3.4{\times}$$7.9{\times}$ and latency by $3.6{\times}$$14.5{\times}$ while maintaining accuracy across math and coding benchmarks. It also surpasses equal-size autoregressive LLMs in tokens-per-second (see Section 5.2).
Figure 1: Left: Block-wise decoding trajectory of the teacher (steps $0\!\to\!N$; diffusion time $t:1\!\to\!0$). Right: The student’s three-objective loss at an intermediate state $y$: (i) distillation from teacher logits on newly unmasked positions, (ii) consistency between $y$ and its block-completion $y^{\star}$, and (iii) masked-denoising (DLM) loss on randomly masked ground-truth text.
2 Background
Here, we briefly review relevant background.
2.1 Diffusion Language Models
Diffusion models ho2020ddpm have been successfully applied across modalities such as text-conditioned image generation ramesh2022hierarchical, audio synthesis chen2021wavegrad, and video generation esser2023sceneguided. In language modeling, diffusion language models (DLMs) generate fixed-length sequences via iterative refinement: at each step, multiple token positions are updated in parallel and finalized gong2023diffuseq; li2022diffusionlm. Early approaches operated in the continuous embedding space gao2024embedding; gong2023diffuseq, but more recent models adopt masked diffusion models (MDMs) that work directly in the discrete token space gong2025scalingdiffusion; nie2025scalingmdm. They begin from a fully masked sequence and iteratively unmask tokens according to a scheduling policy. In this paper, we treat MDMs as the representative class of DLMs, especially Dream ye2025dream7b and LLaDA nie2025large. Architecturally, DLMs commonly employ Transformer backbones with bidirectional (non-causal) attention, permitting full-context interactions nie2025large. Despite their parallel update capability, DLMs are known to be slow for two main reasons. First, full bidirectional attention at every step renders them incompatible with standard KV caching. Second, high-quality generation typically demands a large number of refinement steps comparable to the sequence length, making inference inefficient kim2025beyond. In this work, we propose a fine-tuning-based solution that addresses both limitations.
2.2 Efficient Inference in Diffusion Language Models
Inference in DLMs is an active area of research with significant room for improvement. Among training-free acceleration methods, one line of work exploits approximate caching via block-wise decoding, refreshing static regions periodically ma2025dkvcache; wu2025fastdllm. Others speed up parallel sampling by applying confidence thresholds wu2025fastdllm or by leveraging lightweight autoregressive models as guidance hu2025freecache. In contrast, training-based approaches focus on improving the intrinsic efficiency of DLMs through architectural or objective-level modifications. Some recent works fine-tune models to introduce block-wise causality, enabling standard KV caching at the block level while retaining diffusion-style updates within each block wang2025d2f; wu2025fastdllmv2. Others design objectives that promote faster convergence by encouraging high-confidence token prediction chen2025dparallel. Our method is training-based: we induce block-wise causality and cut refinement steps via fine-tuning.
2.3 Knowledge Distillation in Language Models
Knowledge distillation (KD) hinton2015distilling transfers knowledge from a large, powerful teacher to a smaller, more efficient student. In early LLM work, KD was largely black-box, where students imitated textual outputs from closed-source APIs fu2023specializing. With the rise of open-source LLMs, white-box distillation became feasible, enabling direct supervision from the teacher’s internal representations, such as logits or embeddings agarwal2024onpolicy; gu2023knowledge. Beyond language modeling, distillation has also been successfully applied to diffusion-based image and video generation to cut sampling steps while preserving quality yin2024dmd; yin2025causvid. Our method can similarly be viewed as a form of self-distillation: we distill knowledge from a teacher to a student of identical size and model type, but with different architectural design (a fully bidirectional teacher vs. a block-wise causal student). We adopt a white-box setup with rich intermediate supervision, enabling efficient and stable knowledge transfer. Like deschenaux2025sdtt and xu2025dinfer, we reduce decoding steps by distilling multi-step diffusion trajectories into jumps between non-consecutive denoising states. Our novelty is to instantiate these jumps in a block-causal student, and to train it with consistency-oriented objectives.
2.4 Consistency Models and Language Models
Consistency models song2023consistency were introduced to overcome the slow, iterative sampling process of diffusion models. They learn a consistency property: any intermediate state along a probability flow ODE trajectory can be mapped directly back to its origin (the clean data), enabling near one-step generation. song2023improved. While originally developed for continuous domains such as vision, recent work has explored adapting this idea to discrete language modeling kou2024cllms; xu2025unicms. Rather than explicitly modeling the ODE solver or its continuous trajectory, these methods reinterpret the notion of mapping intermediate states to final solutions within discrete token trajectories. For instance, autoregressive language models, despite lacking an explicit diffusion trajectory, realize this principle via Jacobi-style parallel decoding, treating iterative drafts as intermediate states santilli2023accelerating. Our approach adopts this paradigm in DLMs, where trajectories arise naturally from iterative denoising. Conceptually, our objective can be viewed as an empirical generalization of the original consistency objective defined over continuous ODE trajectories, adapted to discrete, token-level diffusion processes.
3 Preliminary: Inference in Diffusion Language Models
We formalize DLM inference by defining the iterative refinement process, the sampling strategies, and the decoding trajectory that underpins our training objectives.
Iterative Refinement Process.
DLM generation is an iterative refinement over $N$ discrete sampling steps. It gradually transforms a fully masked sequence $\mathbf{x}_{1}$ at time $t=1$  (step 0)  into a clean sequence $\mathbf{x}_{0}$ at $t=0$  (step $N$). Here, $N$ serves as a hyperparameter controlling the trade-off between generation speed and sample quality. We denote the sequence at a discrete step $k$ (where $k=0,1,\dots,N$) as $\mathbf{x}_{t_{k}}$ with $t_{k}=1-\frac{k}{N}$ (so $t_{0}{=}1$, $t_{N}{=}0$). At each step, approximately $L_{g}/N$ tokens are finalized on average, where $L_{g}$ is the target generation length.
This refinement process is modeled probabilistically. Specifically, the transition from a state $\mathbf{x}_{t}$ to a subsequent state $\mathbf{x}_{s}$ $0\leq s<t\leq 1$ is governed by the model $p_{\theta}$’s prediction of the clean sequence $\mathbf{x}_{0}$ given the current noisy sequence $\mathbf{x}_{t}$ and a conditioning prompt $c$, i.e.,
$p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c).$ (1)
Following prior formulations austin2021structured, the reverse-time transition $q_{s|t}(\mathbf{x}_{s}\mid\mathbf{x}_{t})$ is defined for $0\leq s<t\leq 1$ and factorizes across tokens. For each token $i$, the transition probability is defined as:
$q_{s|t}(x_{s}^{i}\mid x_{t}^{i})=\begin{cases}1,&\\ \quad\quad\text{if }x_{t}^{i}\neq\texttt{[MASK]},\;x_{s}^{i}=x_{t}^{i},\\[6.0pt] \frac{s}{t},&\\ \quad\quad\text{if }x_{t}^{i}=\texttt{[MASK]},\;x_{s}^{i}=\texttt{[MASK]},\\[6.0pt] \frac{t-s}{t}\,q_{0|t}(x_{s}^{i}\mid\mathbf{x}_{t},c),&\\ \quad\quad\text{if }x_{t}^{i}=\texttt{[MASK]},\;x_{s}^{i}\neq\texttt{[MASK]}.\end{cases}$ (2)
where $q_{0|t}$ is the predictive distribution induced by $p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c)$. This formulation captures three possibilities: (1) an already-unmasked token remains unchanged, (2) a masked token remains masked, or (3) a masked token is unmasked to a new token.
Sampling Strategies.
To instantiate the probabilistic distribution $q_{s|t}$ into an actual generation procedure, practical deterministic strategies are employed. A common choice is low-confidence remasking, where instead of stochastic sampling, the model greedily selects the tokens to unmask nie2025large. Based on the confidence scores from $p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c)$, the top-$m$ masked tokens with the highest probabilities are revealed, while all others remain masked. In practice, this process is often constrained to operate within blocks of tokens, called block-wise decoding, limiting unmasking to localized regions of the sequence. This prevents unnatural early generation (e.g., premature <endoftext> tokens) nie2025large and, more importantly, enables KV caching for finalized blocks.
Decoding Trajectory.
Given this iterative and probabilistic refinement process, we define the decoding trajectory $\mathcal{T}$ as the sequence of states that the model traverses during inference:
$\mathcal{T}_{\textbf{x}}=\big(\mathbf{x}_{t_{0}},\,\mathbf{x}_{t_{1}},\,\ldots,\,\mathbf{x}_{t_{N}}\big),\quad t_{k}=1-\frac{k}{N}.$ (3)
Each $\mathbf{x}_{t_{k}}$ is a partially refined sequence; the trajectory $\mathcal{T}_{\mathbf{x}}$ records its evolution across steps. Figure 1 (left) visualizes an example. This trajectory serves as the foundation for our consistency-based training below.
4 Methodology
We present Consistency Diffusion Language Models (CDLM), a training scheme that accelerates DLM inference without sacrificing quality. We first describe how we collect teacher decoding trajectories and store compact hidden states for logit reconstruction, then introduce the distillation, consistency, and DLM objectives, and finally detail our confidence-thresholded block-wise inference.
Figure 2: Left: Teacher DLM with full bidirectional attention, attending to the entire context. Right: Student DLM with a block-wise causal mask, attending to the prompt, previously completed blocks, and the current decoding block.
Algorithm 1 Trajectory Collection for Training CDLM
[…]
2:Output: dataset $\mathcal{D}$ of quadruples $(x,\hat{\mathbf{y}},\mathcal{T}_{x},\mathbf{H}_{x})$
3: Initialize $\mathcal{D}\leftarrow\emptyset$
4:repeat
5:  Sample $(x,\hat{\mathbf{y}})\sim\mathcal{O}$; build input ids $\mathbf{x}$ with the answer span masked
6:  for each temperature $\tau\in\mathcal{S}_{\tau}$ do
7:   $\mathcal{T}\leftarrow[\,]$;  $\mathbf{H}\leftarrow\mathbf{0}\in\mathbb{R}^{L_{g}\times d}$
8:   for blocks $b=1,\dots,\lceil L_{g}/B\rceil$ do
9:    for steps $s=1,\dots,B$ do
10:     Run $p_{\theta}(\mathbf{x})$; obtain logits $\ell$ and last hidden $\mathbf{h}$
11:     Sample $\mathbf{y}$ and compute confidence $\mathbf{c}$
12:     Finalize top-1 token $y_{i}$ in the current block; write $\mathbf{h}_{i}\!\to\!\mathbf{H}$; append $\mathbf{x}\!\to\!\mathcal{T}$
13:    end for
14:   end for
15:   Append $(x,\hat{\mathbf{y}},\mathcal{T},\mathbf{H})$ to $\mathcal{D}$
16:  end for
17:until $|processed\_pairs|\geq n$
4.1 Trajectory Collection for Diffusion Language Models
Data Generation
Let $p_{\theta}$ denote the teacher DLM and $q_{\phi}(\cdot\mid x)$ the student DLM initialized from the teacher’s weights. The teacher uses a bidirectional attention mask as in standard DLMs, whereas the student is trained with a block-wise causal mask, as illustrated in Figure 2. To enable consistency training, we collect trajectories offline by running the teacher on domain-specific prompts. For each prompt $x$, we denote the token-level trajectory as $\mathcal{T}_{x}$ and the corresponding hidden-state buffer by $\mathbf{H}_{x}\in\mathbb{R}^{L_{g}\times d}$; the resulting dataset is $\mathcal{D}=\{(x,\hat{\mathbf{y}},\mathcal{T}_{x},\mathbf{H}_{x})\}$ with ground-truth text $\hat{\mathbf{y}}$ included. The overall procedure is summarized in Algorithm 1.
Prior work reports that block-wise decoding is effective for instruction-tuned models and attains high quality when the number of steps matches the generation length nie2025large. Accordingly, we adopt block-wise decoding with the number of sampling steps equal to the generation length ($N{=}L_{g}$) and finalize exactly one token per step within the current block (the highest-confidence token $y_{i}$). This configuration places the teacher at its most performant operating point and yields higher-quality trajectories. For white-box distillation, we store not only $\mathcal{T}_{x}$ but also the states required to reconstruct logits. Concretely, we record the last hidden states at the moments when tokens are finalized.
Data Augmentation
Effective training benefits from trajectories that are both diverse and reliable. Unlike autoregressive models, where temperature only influences token selection, in DLMs, the temperature also changes the order in which tokens are revealed gong2025diffucoder which allows further diversity in the generated data. We therefore augment each prompt with multiple trajectories generated at different temperatures. Additional details on dataset construction are provided in Appendix A.1.
Algorithm 2 Training Algorithm for CDLM
1:Input: dataset $\mathcal{D}=\{(x,\hat{\mathbf{y}},\mathcal{T}_{x},\mathbf{H}_{x})\}$, block size $B$, generation length $L_{g}$, weights $w_{\text{cons}},w_{\text{distill}}$, $w_{\text{dlm}}$, student DLM $q_{\phi}(\cdot\mid x)$ with block-wise causal attention, total steps $N{=}L_{g}$
2:Output: updated student parameters $\phi$
3:repeat
4:  Sample $(x,\hat{\mathbf{y}},\mathcal{T}_{x},\mathbf{H}_{x})\sim\mathcal{D}$
5:  Sample $t_{\text{start}}$; set $t_{\text{end}}\leftarrow\min\!\left(N,\,\left\lceil\tfrac{t_{\text{start}}}{B}\right\rceil B\right)$
6:  Retrieve states at $t_{\text{start}}$ and $t_{\text{end}}$ from $\mathcal{T}_{x}$
7:  Compute $\mathcal{L}_{\text{Distillation}}$ (Eq. (4)) using $\mathbf{H}_{x}$
8:  Compute $\mathcal{L}_{\text{Consistency}}$ (Eq. (5))
9:  Compute $\mathcal{L}_{\text{DLM}}$ (Eq. (6)) using ground-truth text $\hat{\mathbf{y}}$
10:  $\mathcal{L}\leftarrow w_{\text{distill}}\mathcal{L}_{\text{Distillation}}+w_{\text{cons}}\mathcal{L}_{\text{Consistency}}+w_{\text{dlm}}\mathcal{L}_{\text{DLM}}$
11:  Update $\phi$ with $\mathcal{L}$
12:until max epochs
4.2 Training
We jointly minimize three objectives to train CDLMs: (i) a Distillation loss that transfers the teacher’s multi-token finalization signal to the block-wise causal student; (ii) a Consistency loss that enforces within-block temporal consistency between two states; and (iii) a DLM loss that preserves the model’s masked-token prediction capability. The detailed training procedure is depicted in Algorithm 2.
Notation
Given the dataset $\mathcal{D}=\{(x,\hat{\textbf{y}},\mathcal{T}_{x},\mathbf{H}_{x})\}$, we first sample a prompt $x$ and its trajectory $\mathcal{T}_{x}$, then sample a state $y\in\mathcal{T}_{x}$ and its block-completion state $y^{\star}$. Here, $y^{\star}$ denotes the state obtained by fully unmasking the active block of $y$ (thus $y$ and $y^{\star}$ are at most $B$ steps apart). The right panel of Figure 1 visualizes this setup. We denote $\mathcal{U}_{y}=\{i\mid y_{i}=\texttt{[MASK]},\,y_{i}^{\star}\neq\texttt{[MASK]}\}$ as the newly unmasked token indices between $y$ and $y^{\star}$, and $\mathcal{S}_{y}=\{i\mid y_{i}=\texttt{[MASK]},\,y_{i}^{\star}=\texttt{[MASK]}\}$ as the still-masked token indices at $y^{\star}$.
Distillation Loss
We use the student’s predictions at state $y$ together with the teacher’s hidden-state buffer $\mathbf{H}_{x}$. For each $i\in\mathcal{U}_{y}$, we reconstruct the teacher’s distribution by applying $\ell^{(T)}_{i}=\texttt{lm\_head}(\mathbf{h}_{x,i})$ and $p^{(T)}_{i}=\operatorname{softmax}(\ell^{(T)}_{i})$ to the stored hidden state $\mathbf{h}_{x,i}$. With these teacher distributions, we define the distillation objective as
$\mathcal{L}_{\text{Distillation}}$ $=\mathbb{E}_{(x,\mathcal{T}_{x},\mathbf{H}_{x})\sim\mathcal{D}}\,\mathbb{E}_{y\sim\mathcal{T}_{x}}\Bigg[$ (4)
$\qquad\frac{1}{|\mathcal{U}_{y}|}\sum_{i\in\mathcal{U}_{y}}D_{\mathrm{KL}}\!\Big(p^{(T)}_{i}\,\big\|\,q_{\phi}(\,\cdot\mid y,x)_{i}\Big)\Bigg],$
which applies forward KL divergence on the positions newly unmasked between $y$ and $y^{\star}$. This serves as the main anchor that transfers the bidirectional teacher’s guidance to the block-wise student, effectively providing supervision for multi-token finalization within each block.
Consistency Loss
We compare the student’s predictive distributions at state $y$ and block-completion state $y^{\star}$, enforcing agreement on still-masked positions. Concretely, we minimize the following objective:
$\mathcal{L}_{\text{Consistency}}$ $=\mathbb{E}_{(x,\mathcal{T}_{x})\sim\mathcal{D}}\,\mathbb{E}_{y\sim\mathcal{T}_{x}}\Bigg[$ (5)
$\hskip-15.00002pt\frac{1}{|\mathcal{S}_{y}|}\sum_{i\in\mathcal{S}_{y}}D_{\mathrm{KL}}\!\Big(q_{\phi^{-}}(\,\cdot\mid y^{\star},x)_{i}\;\big\|\;q_{\phi}(\,\cdot\mid y,x)_{i}\Big)\Bigg].$
Here, $q_{\phi^{-}}$ denotes a stop-gradient target (detached from backpropagation) for stable training song2023consistency, and $D_{\mathrm{KL}}$ is the forward KL divergence between two distributions. Intuitively, Eq. (5) aligns the student at a less-informed state with itself at a more-informed state, but only over $\mathcal{S}_{y}$ where predictions are well-defined. Since DLMs are trained to predict only [MASK] tokens, restricting the loss to still-masked indices avoids ill-defined supervision and helps the model jump several steps in a stable, consistent manner.
DLM loss
We include an auxiliary loss, identical to the masked denoising objective used in standard DLM pre-training nie2025large; ye2025dream7b, to preserve the model’s mask prediction capability. For a prompt $x$ and its ground-truth response $\hat{\mathbf{y}}$, we randomly sample a masking ratio $t\sim\mathcal{U}[0,1]$ and independently mask each token with a probability $t$ to obtain the masked sequence $\hat{\mathbf{y}}_{t}$. The resulting objective is
$\mathcal{L}_{\text{DLM}}$ $=-\mathbb{E}_{(x,\hat{\mathbf{y}})\sim\mathcal{D}}\,\mathbb{E}_{t}\Bigg[$ (6)
$\qquad\frac{1}{t}\sum_{i=1}^{L_{g}}\mathbf{1}\!\big[\hat{y}_{t,i}=\texttt{[MASK]}\big]\,\log q_{\phi}\big(\hat{y}_{i}\mid\hat{\mathbf{y}}_{t},x\big)\Bigg].$
Here, $\mathbf{1}[\cdot]$ denotes the indicator function.
Consequently, the total training objective is
$\mathcal{L}(\phi)=w_{\text{distill}}\,\mathcal{L}_{\text{Distillation}}+w_{\text{cons}}\,\mathcal{L}_{\text{Consistency}}+w_{\text{dlm}}\,\mathcal{L}_{\text{DLM}}.$ (7)
4.3 Inference
At inference time, the student decodes block-wise under a block-causal mask (Figure 2), reusing the KV cache for the prompt and all previously finalized blocks. Within each block, we apply confidence-thresholded parallel finalization: at each step, among the masked positions of the current block, we reveal every token whose confidence exceeds a threshold $\tau_{\text{conf}}$, following Fast-dLLM wu2025fastdllm. Larger $\tau_{\text{conf}}$ yields more conservative decoding (typically higher quality but slower convergence), whereas smaller $\tau_{\text{conf}}$ is more aggressive (faster convergence) at the cost of potential accuracy loss. We also adopt early stopping: decoding terminates once an <endoftext> token is produced within the current block, analogous to AR decoding. We intentionally avoid additional heuristics such as inter-block parallelism wang2025d2f, which introduce extra hyperparameters whose optimal values are task- and domain-dependent.
Table 1: Evaluation results for Dream-7B-Instruct. Arrows in headers indicate whether higher ($\uparrow$) or lower ($\downarrow$) is better. Notes: Par. = parallel decoding; D.C. = dual-cache KV.
Benchmark Method TPS $\uparrow$ Latency (s) $\downarrow$ Total Steps $\downarrow$ Gen. Length Score $\uparrow$
GSM8K -CoT (8-shot) Dream-7B-Instruct 4.1 (×1.0) 23.5 (×1.0) 256.0 (×1.0) 95.1 79.1
dLLM-Cache 6.9 (×1.7) 12.6 (×1.9) 256.0 (×1.0) 87.5 75.2
Fast-dLLM (Par.) 19.2 (×4.7) 5.0 (×4.7) 53.7 (×4.8) 96.0 79.9
Fast-dLLM (Par.+D.C.) 36.6 (×8.9) 2.5 (×9.4) 60.8 (×4.2) 90.3 77.3
CDLM–Dream (ours) 51.7 (×12.6) 2.1 (×11.2) 44.1 (×5.8) 107.4 78.8
HumanEval -Instruct (0-shot) Dream-7B-Instruct 15.4 (×1.0) 13.4 (×1.0) 256.0 (×1.0) 206.7 48.2
dLLM-Cache 9.4 (×0.6) 12.0 (×1.1) 256.0 (×1.0) 113.6 43.9
Fast-dLLM (Par.) 66.4 (×4.3) 3.2 (×4.2) 61.6 (×4.2) 210.2 45.7
Fast-dLLM (Par.+D.C.) 79.9 (×5.2) 2.5 (×5.4) 71.6 (×3.6) 200.9 46.3
CDLM–Dream (ours) 43.3 (×2.8) 2.2 (×6.1) 49.6 (×5.2) 96.9 50.0
MATH (4-shot) Dream-7B-Instruct 8.3 (×1.0) 21.9 (×1.0) 256.0 (×1.0) 182.2 38.0
dLLM-Cache 11.3 (×1.4) 13.0 (×1.7) 256.0 (×1.0) 146.9 35.8
Fast-dLLM (Par.) 23.9 (×2.9) 7.6 (×2.9) 87.1 (×2.9) 181.0 37.8
Fast-dLLM (Par.+D.C.) 48.8 (×5.9) 3.7 (×5.9) 98.2 (×2.6) 179.8 37.4
CDLM–Dream (ours) 53.8 (×6.5) 2.9 (×7.6) 63.2 (×4.1) 158.4 32.4
MBPP -Instruct (0-shot) Dream-7B-Instruct 2.3 (×1.0) 21.7 (×1.0) 256.0 (×1.0) 49.5 51.8
dLLM-Cache 5.1 (×2.2) 11.3 (×1.9) 256.0 (×1.0) 57.3 56.0
Fast-dLLM (Par.) 15.5 (×6.7) 3.1 (×7.0) 35.3 (×7.3) 47.3 55.6
Fast-dLLM (Par.+D.C.) 25.4 (×11.0) 1.9 (×11.4) 43.6 (×5.9) 47.1 53.4
CDLM–Dream (ours) 48.1 (×20.9) 1.5 (×14.5) 33.2 (×7.7) 74.3 53.0
Table 2: Evaluation results for LLaDA-8B-Instruct. Arrows in headers indicate whether higher ($\uparrow$) or lower ($\downarrow$) is better. We report HumanEval (not HumanEval-Instruct), as the latter lowers overall baseline scores. Notes: Par. = parallel decoding; D.C. = dual-cache KV.
Benchmark Method TPS $\uparrow$ Latency (s) $\downarrow$ Total Steps $\downarrow$ Gen. Length Score $\uparrow$
GSM8K (4-shot) LLaDA-8B-Instruct 8.2 (×1.0) 28.3 (×1.0) 256.0 (×1.0) 232.1 77.1
dLLM-Cache 18.7 (×2.3) 12.3 (×2.3) 256.0 (×1.0) 229.6 78.5
Fast-dLLM (Par.) 26.7 (×3.3) 8.7 (×3.3) 77.5 (×3.3) 231.3 78.6
Fast-dLLM (Par.+D.C.) 54.6 (×6.7) 4.2 (×6.7) 85.0 (×3.0) 230.4 76.5
CDLM–LLaDA (ours) 54.3 (×6.6) 3.3 (×8.6) 57.7 (×4.4) 177.3 73.9
HumanEval (0-shot) LLaDA-8B-Instruct 7.4 (×1.0) 11.3 (×1.0) 256.0 (×1.0) 83.9 37.8
dLLM-Cache 12.6 (×1.7) 10.8 (×1.0) 256.0 (×1.0) 135.9 39.0
Fast-dLLM (Par.) 17.9 (×2.4) 4.6 (×2.5) 100.3 (×2.6) 83.1 36.6
Fast-dLLM (Par.+D.C.) 18.9 (×2.6) 4.2 (×2.7) 97.7 (×2.6) 80.2 36.0
CDLM–LLaDA (ours) 50.9 (×6.9) 1.9 (×5.9) 32.3 (×7.9) 95.1 40.2
MATH (4-shot) LLaDA-8B-Instruct 8.8 (×1.0) 25.7 (×1.0) 256.0 (×1.0) 226.4 24.1
dLLM-Cache 22.7 (×2.6) 10.9 (×2.4) 256.0 (×1.0) 248.5 33.3
Fast-dLLM (Par.) 25.1 (×2.9) 9.9 (×2.6) 96.9 (×2.6) 248.6 33.4
Fast-dLLM (Par.+D.C.) 49.7 (×5.7) 5.0 (×5.1) 107.0 (×2.4) 248.2 32.5
CDLM–LLaDA (ours) 50.2 (×5.7) 4.2 (×6.1) 75.3 (×3.4) 210.3 28.3
MBPP -Instruct (0-shot) LLaDA-8B-Instruct 17.7 (×1.0) 11.4 (×1.0) 256.0 (×1.0) 201.0 40.8
dLLM-Cache 18.5 (×1.0) 10.8 (×1.1) 256.0 (×1.0) 199.4 40.8
Fast-dLLM (Par.) 21.2 (×1.2) 6.2 (×1.8) 59.1 (×4.3) 131.8 42.0
Fast-dLLM (Par.+D.C.) 40.1 (×2.3) 3.4 (×3.4) 66.9 (×3.8) 134.7 35.0
CDLM–LLaDA (ours) 60.6 (×3.4) 3.2 (×3.6) 58.0 (×4.4) 195.3 38.4
5 Experiments
5.1 Experimental Setup
This subsection summarizes datasets, CDLM training, evaluation protocol, baselines, hardware, and metrics.
Target Models and Datasets
We evaluate two representative open-source DLMs: Dream-7B-Instruct ye2025dream7b and LLaDA-8B-Instruct nie2025large. Our prompts are derived from Bespoke-Stratos-17k bespoke_stratos. Specifically, we use two publicly available datasets from the Hugging Face Hub lansechen_easy_2025; lansechen_hard_2025, retaining items whose prompt length is $\leq 512$ tokens. For LLaDA, we further augment its corpus with 7.57.5k additional math-style prompts sourced from the DParallel dataset chen2025dparallel. For each prompt, we generate teacher trajectories at temperatures $\{0.0,\,0.5\}$ (greedy / non-greedy), yielding 1515k generations for Dream and 3030k for LLaDA. Generation uses the bidirectional, block-wise teacher with generation length $L_{g}{=}256$, block size B=32B{=}32, and total steps N=256N{=}256. Additional details are provided in Appendix A.1.
Training Details
[…]
Benchmarks.
Following established conventions wang2025d2f; wu2025fastdllm, we evaluate mathematical reasoning and code generation on GSM8K, GSM8K-CoT cobbe2021gsm8k, MATH hendrycks2021math, HumanEval chen2021humaneval, and MBPP austin2021mbpp. We use the lm_eval framework eval-harness and the Instruct task variants for the coding benchmarks. We report exact match scores on GSM8K, math-verify scores on MATH, and pass@1 results on HumanEval, HumanEval-Instruct, and MBPP-Instruct after standard post-processing. Coding tasks are evaluated zero-shot, and few-shot settings for math follow the original Dream and LLaDA papers. All efficiency measurements are taken on $4\times$ NVIDIA A100 (80 GB) GPUs with batch size $1$ under data parallelism. Latency, total steps, and generation length are reported as per-sample averages over the evaluation set.
Baselines.
As references, we evaluate the original DLMs under their official inference settings. For LLaDA, we enable block-wise decoding. Dream does not support block-wise decoding, so we extend the baseline accordingly. Because our work targets two orthogonal axes of inference acceleration of (1) KV caching and (2) step reduction, we include three inference-time baselines that isolate or combine these factors: (i) dLLM-Cache liu2025dllmcache, which focuses on (1) via adaptive feature caching; (ii) Fast-dLLM (Parallel) wu2025fastdllm, which targets (2) via confidence thresholding; and (iii) Fast-dLLM (Parallel + Dual Cache) wu2025fastdllm, which combines (1) + (2) via approximate dual-cache KV caching. We omit D2F wang2025d2f here because it is trained for $L_{g}{=}512$ and reports on Dream-7B-Base rather than Dream-7B-Instruct. For autoregressive comparisons, we evaluate Qwen2.5-7B-Instruct qwen2024techreport alongside Dream and Llama-3.1-8B-Instruct dubey2024llama alongside LLaDA. For fair comparison, we use a fixed generation length of $L_{g}{=}256$, greedy decoding (temperature $0.0$), a block size of B=32B{=}32 for block-wise decoding, and a token-confidence threshold of $\tau_{\text{conf}}{=}0.9$. Further details are in Appendix A.3.
5.2 Main Results
We compare CDLM against vanilla DLMs, accelerated DLM variants, and autoregressive baselines, and summarize the quantitative results.
5.2.1 Results on CDLM–Dream
Table 1 summarizes throughput, latency, step counts, generation length, and accuracy for Dream-7B-Instruct and its acceleration baselines.
Steps and scores.
[…]
We attribute the drop on MATH to two factors. First, the small training mixture ($\sim$7.5k prompts) limits exposure to the advanced problem-solving skills required by MATH. Since the mixture is drawn from filtered subsets of Bespoke-Stratos-17k that largely consist of problems Qwen2.5 can already solve, the effective difficulty range is likely narrower than that of MATH. Second, CDLM is a fine-tuning-based method trained with teacher trajectories capped at 256 generation tokens. While sufficient for other benchmarks, this shorter thinking budget may be inadequate for complex multi-step reasoning. A natural next step would be to train with a 512-token budget, as D2F wang2025d2f does.
Latency.
[…]
Throughput (TPS).
CDLM–Dream achieves the highest TPS on three out of four tasks, outperforming all baselines. On MBPP-Instruct, it jumps from 2.3 to 48.1 tokens/s (a $20.9\times$ increase over the naive model). The one exception is HumanEval-Instruct, where our TPS is lower than Fast-dLLM baselines because our effective generation length is much shorter (about 97 tokens vs. $\sim$200 for the baselines). This can depress our measured TPS and inflate baseline TPS. Crucially, despite shorter generations, our pass@1 is higher, indicating that CDLM–Dream emits fewer redundant tokens while maintaining or improving solution quality.
5.2.2 Results on CDLM–LLaDA
Table 2 summarizes throughput, latency, step counts, generation length, and accuracy for LLaDA-8B-Instruct and its acceleration baselines.
Steps and scores.
[…]
Latency and Throughput (TPS).
[…]
Figure 3: Throughput vs. AR (Dream). Tokens-per-second on GSM8K-CoT and MBPP-Instruct for Dream-7B-Instruct (naive), Qwen2.5-7B-Instruct (AR), and CDLM–Dream.
Figure 4: Throughput vs. AR (LLaDA). Tokens-per-second on GSM8K and HumanEval for LLaDA-8B-Instruct (naive), Llama3.1-8B-Instruct (AR), and CDLM–LLaDA.
5.2.3 Comparison with Autoregressive Models
Throughput.
In terms of throughput, CDLM improves naive DLMs by $3\times$$21\times$ and also surpasses equal-size autoregressive (AR) baselines. For Dream, we compare against Qwen2.5-7B-Instruct (Figure 3), where CDLM–Dream achieves $1.21\times$ higher throughput on GSM8K-CoT and $1.12\times$ on MBPP-Instruct. For LLaDA, we compare against Llama-3.1-8B-Instruct (Figure 4), where CDLM–LLaDA reaches $1.27\times$ the throughput on GSM8K and $4.17\times$ on HumanEval. Throughput here reflects both per-step cost and the number of refinement steps. A CDLM refinement step is more expensive than an AR step due to matrix-matrix multiplications for decoding. However, CDLM can finalize multiple tokens per iteration, reducing total steps and increasing effective throughput. For example, from Table 1, CDLM–Dream produces on average $107.4/44.1\approx 2.4$ tokens per step on GSM8K-CoT and $74.3/33.2\approx 2.2$ tokens per step on MBPP-Instruct.
Accuracy.
Although throughput is our focus, accuracy may be lower than that of AR models because CDLMs remain bounded by the strength of their DLM backbones. For CDLM–Dream versus its AR baseline, the accuracies are $78.8$ vs. $73.8$ on GSM8K-CoT and $53.0$ vs. $81.7$ on MBPP-Instruct (CDLM vs. AR). For CDLM–LLaDA versus its AR baseline, the respective accuracies are $73.9$ vs. $80.3$ on GSM8K and $40.2$ vs. $60.4$ on HumanEval (CDLM vs. AR). As stronger DLM backbones emerge, applying our consistency-based fine-tuning should yield further throughput gains with improved quality.
5.3 Ablation Studies
We analyze how training and inference choices affect efficiency and accuracy.
Table 3: Ablation of loss weights. We vary $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ and report score (top, black) and steps to convergence (bottom, blue) on GSM8K (4-shot) and HumanEval-Instruct (0-shot) for CDLM–Dream.
$\boldsymbol{w_{\text{distill}}}$ $\boldsymbol{w_{\text{cons}}}$ $\boldsymbol{w_{\text{dlm}}}$ GSM8K (4-shot) HumanEval-Inst. (0-shot)
1.0 0.01 73.2 (46.7) 42.7 (61.0)
1.0 0.01 6.9 (100.6) 0.0 (124.3)
1.0 1.0 0.01 74.1 (49.4) 42.7 (49.4)
1.0 1.0 73.3 (46.2) 48.2 (51.5)
1.0 0.1 0.01 75.1 (48.0) 45.7 (59.9)
1.0 0.1 73.7 (48.1) 50.6 (69.0)
5.3.1 Loss-Weight Composition
Table 3 varies the loss weights $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ to study the trade-off between convergence speed and final accuracy. We train the Dream model for four epochs with a constant learning rate 2e-5, then evaluate on GSM8K (4-shot) and HumanEval-Instruct (0-shot).
Distillation vs. consistency in isolation.
With a small auxiliary DLM loss ($w_{\text{dlm}}{=}0.01$), the distillation-only setting (row 1) acts as a strong anchor: it converges quickly (about 46.7 and 61.0 steps for a generation length of 256), with a slight degradation in score. In contrast, consistency-only (row 2) collapses: optimizing for self-agreement without teacher supervision causes failure.
Coupling yields synergy.
[…]
Role of the auxiliary DLM loss.
Removing the ground-truth masked-denoising term (rows 4 and 6) raises HumanEval but lowers GSM8K, indicating a loss of math reasoning ability. A small auxiliary weight helps preserve this ability, and we find that general coding performance can be recovered with longer training.
Based on these trends, we fix $w_{\text{distill}}{=}1.0$ and empirically choose $w_{\text{cons}}{=}0.5$ as a good balance between speed and quality. For the auxiliary DLM loss, we use $w_{\text{dlm}}{=}0.01$ for Dream and $w_{\text{dlm}}{=}0.1$ for LLaDA, since the DLM loss on LLaDA has a smaller absolute scale.
Table 4: Ablation of refinement steps. We report GSM8K results after naively truncating baseline DLM step counts to approximately match the step budgets used by CDLM–Dream and CDLM–LLaDA.
Method Latency (s) $\downarrow$ Steps $\downarrow$ Score $\uparrow$
Dream-7B-Instruct 4.4 48 41.8
CDLM–Dream (ours) 2.1 44.1 78.8
LLaDA-8B-Instruct 6.0 56 60.3
CDLM–LLaDA (ours) 3.3 57.7 73.9
5.3.2 Effective Step Reduction
We evaluate a naive step-truncation baseline by forcing the teacher DLMs to use a similar number of refinement steps as our CDLMs (Table 4). Because decoding is block-wise, the step count must be a multiple of the block factor ($L_{g}/B{=}8$). For Dream, we set the step budget to $48$ and measure GSM8K-CoT (8-shot) as in Table 1; for LLaDA, we set it to $56$ and measure GSM8K (4-shot) as in Table 2. Naively truncating the step budget, in other words, forcing a non-retrained DLM to finalize multiple tokens per iteration, leads to marked accuracy degradation. By contrast, CDLM maintains quality while operating with substantially fewer steps, underscoring that consistency training is necessary for stable multi-token refinement. Furthermore, with KV caching, our approach reduces latency by roughly half relative to the naive DLM at comparable iteration counts.
Table 5: Ablation of the confidence threshold. We sweep the token-level threshold $\tau_{\text{conf}}$ for CDLM–Dream on GSM8K-CoT (8-shot) and HumanEval-Instruct (0-shot).
Benchmark $\boldsymbol{\tau_{\text{conf}}}$ TPS $\uparrow$ Latency (s) $\downarrow$ Score $\uparrow$
GSM8K
-CoT
(8-shot)
0.95 42.7 2.5 78.8
0.90 51.7 2.1 78.8
0.85 57.7 1.9 78.4
HumanEval
-Instruct
(0-shot)
0.95 34.4 2.8 51.2
0.90 43.3 2.2 50.0
0.85 47.3 2.0 48.2
5.3.3 Token-Level Confidence Threshold
Table 5 sweeps the token-level confidence threshold $\tau_{\text{conf}}\!\in\!\{0.85,0.90,0.95\}$ to characterize conservative-aggressive decoding behavior. Increasing $\tau_{\text{conf}}$ makes finalization more conservative, tokens are only accepted when the model is highly confident, so throughput (TPS) decreases and per-sample latency increases. Decreasing $\tau_{\text{conf}}$ makes decoding more aggressive and has the opposite effect. This monotonic speed trend holds on both GSM8K-CoT and HumanEval-Instruct. Concretely, on GSM8K-CoT, TPS/latency moves from $42.7/2.5$ s at $\tau{=}0.95$ to $57.7/1.9$ s at $\tau{=}0.85$; on HumanEval-Instruct, from $34.4/2.8$ s to $47.3/2.0$ s.
Accuracy shows a similar monotonic trend in $\tau_{\text{conf}}$. On GSM8K-CoT, scores are essentially flat at $78.8$ for $\tau{=}0.95$ and $0.90$, with a slight drop at $\tau{=}0.85$ ($78.4$). Similarly, on HumanEval-Instruct, the best score appears at the most conservative setting ($51.2$ at $\tau{=}0.95$), 3 points higher than the most aggressive setting ($\tau{=}0.85$). These results suggest that raising $\tau_{\text{conf}}$ trades speed for quality, but the gain is not linear. In practice, we find $\tau_{\text{conf}}{=}0.90$ to be a robust default that balances speed and quality across tasks.
6 Discussion
Here, we discuss some complementary and future avenues for our work, as well as its limitations.
Inference-only accelerations.
Our method complements and can be combined with various inference-only (non-trainable) acceleration techniques. At inference time, we use only block-wise KV caching and confidence-based multi-token finalization, but more sophisticated mechanisms can be layered on top. For example, one can incorporate D2F’s inter-block parallelism to further reduce wall-clock latency wang2025d2f.
DLM as a drafter.
Speculative decoding employs a smaller draft model to propose token bursts that a larger verifier accepts or rejects in parallel leviathan2023fast. A naive diffusion language model (DLM) drafter yields limited benefit because standard DLMs require many refinement steps to produce coherent text. In contrast, a consistency-trained DLM (CDLM) can generate reasonable drafts in far fewer steps. Pairing a CDLM drafter with an autoregressive (AR) verifier (cf. christopher2024specdiff) is promising: the CDLM proposes multi-token candidates quickly, and the AR model supplies strong final validation.
Limitations.
Training currently relies on offline, static trajectories. Despite careful dataset selection, the student may overfit to the teacher’s prior because supervision does not adapt during learning. For LLaDA, we observe that augmenting the training set with additional math-style prompts improves math performance by 2–3 percentage points while preserving other abilities, suggesting that scale and domain coverage are critical factors. Scaling beyond our current 15k-prompt corpus and broadening it to more diverse domains are straightforward directions for future work. Furthermore, CDLM’s performance is bounded by the teacher: a bidirectional DLM distilled into a block-causal student cannot exceed the teacher’s knowledge. Distilling from stronger DLM or AR teachers is a natural way to lift this ceiling. Additional discussion appears in Appendix B.
7 Conclusion
We presented CDLM, a training-based acceleration scheme that brings consistency modeling to DLMs. By enforcing within-block temporal consistency and fine-tuning a block-wise causal student, CDLM reduces refinement steps and enables natural KV caching. Across math and coding tasks, it yields faster inference, fewer steps, lower latency, and higher throughput, while maintaining competitive accuracy. Promising directions include scaling the distillation corpus, broadening domain coverage, and distilling from stronger teachers.
Acknowledgements
We acknowledge gracious support from the FuriosaAI, Intel, Apple, NVIDIA, Macronix, and Mozilla team. Furthermore, we appreciate support from Google Cloud, the Google TRC team and Prof. David Patterson. Prof. Keutzer’s lab is sponsored by the Intel corporation, UC Berkeley oneAPI Center of Excellence, Intel VLAB team, as well as funding through BDD and BAIR. We also acknowledge support by the Director, Office of Science, Office of Advanced Scientific Computing Research, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. DOE SciGPT grant. Minseo also gratefully acknowledges the SNUCSE GPU Service 3.0 program and Bacchus for providing A100 GPU resources that supported this research. Our conclusions do not necessarily reflect the position or the policy of our sponsors, and no official endorsement should be inferred.
[…]
Figure 6: Teacher trajectory and hidden-state buffer. Left: Block-wise decoding trajectory. Right: As each token is finalized, the teacher’s last hidden state at that position is written to a fixed-length buffer of size $L_{g}$; arrows indicate write indices across steps (toy example with B=3B{=}3, $L_{g}{=}6$).
Appendix A Additional Experimental Details
A.1 Dataset and Preprocessing
Sources.
We use two datasets lansechen_easy_2025; lansechen_hard_2025, which are post-processed versions of Bespoke-Stratos 17k bespoke_stratos. We choose them because (i) D2F wang2025d2f, a representative fine-tuning-based DLM acceleration method, uses the same source, and (ii) they are filtered to a maximum length of 600 tokens and predominantly contain math word-style reasoning problems. For both Dream and LLaDA, we treat Qwen2.5-7B’s responses as ground-truth answers.
Dream is trained only on this Bespoke-derived subset. To strengthen math performance for LLaDA, we further augment its corpus with an additional 7.57.5k math-style prompts sampled from the DParallel dataset chen2025dparallel, matching the number of Bespoke-derived prompts. The released DParallel corpus does not include Qwen2.5-7B outputs, so we generate reference answers with Qwen2.5-7B for these prompts, yielding an LLaDA training set whose size is approximately $2\times$ that of Dream. After trying multiple datasets, we empirically observe that dataset selection is crucial; overly rigid formats (e.g., multiple-choice math) tend to hurt the model’s ability to generalize.
Temperature augmentation.
In DLMs, temperature influences not only token choices but also the order in which tokens are revealed gong2025diffucoder. We therefore generate multiple trajectories per prompt at different temperatures. However, as shown in Figure 5, $\tau{=}1.0$ often destabilizes the reasoning chain and can yield incorrect conclusions. Consequently, we restrict trajectory collection to $\tau\in\{0.0,\,0.5\}$.
Hidden-state buffer.
Beyond token trajectories, we store the teacher’s decisions in a compact hidden-state buffer. Whenever a token is finalized, we record the teacher’s last-layer hidden state at that position; over a generation of length $L_{g}$, this yields a buffer of shape $d\times L_{g}$ per example (see Figure 6). Directly saving logits is impractical because their dimensionality is $|V|$ (vocabulary size); using hidden states of dimension $d$ instead gives roughly a $30\times$ reduction in storage. The trajectory datasets fit on a single server: each shard is $25\text{-}30\,\mathrm{GiB}$ for $15\mathrm{k}$ samples.
A.2 Training
Training Configurations.
Table 6: Training configuration for CDLM–Dream-7B-Instruct.
Learning rate $2\times 10^{-5}$
Scheduler (warmup) constant (5%)
Epochs (best) 16 (best: 12)
Effective batch size 64
Optimizer AdamW
LoRA rank / $\alpha$ 32 / 32
LoRA targets q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj
Prompt / generation length 512 / 256
Loss weights $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ (1.0,  0.5,  0.01)
Table 7: Training configuration for CDLM–LLaDA-8B-Instruct.
Learning rate $1\times 10^{-5}$
Scheduler (warmup) constant (5%)
Epochs (best) 16 (best: 12)
Effective batch size 64
Optimizer AdamW
LoRA rank / $\alpha$ 64 / 64
LoRA targets q_proj, k_proj, v_proj, attn_out, up_proj, ff_proj, ff_out
Prompt / generation length 512 / 256
Loss weights $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ (1.0,  0.5,   0.1)
Full configurations are listed in Tables 6 and 7. Training ran on $4\times$ NVIDIA A100 (80 GB) or $8\times$ NVIDIA RTX A6000 (48 GB) GPUs; each epoch took roughly $30$ minutes for 1515k trajectories. We used per-device batch sizes of 1122 with gradient accumulation to keep the effective batch size fixed at $64$.
Loss formulations.
Empirically, the forward KL divergence yielded more stable, better-calibrated training dynamics and more monotonic convergence than the reverse KL divergence for distribution matching. We also found that distillation in logit space outperformed embedding-space distillation using mean squared error (MSE).
A.3 Evaluation
Hardware and parallelism.
[…]
Tooling and post-processing.
We use lm_eval (EleutherAI’s lm-eval-harness) for dataset loading, prompting, and scoring. We apply the harness’s default post-processors. Decoded text is truncated at any task-specific stop_sequences prior to scoring. For HumanEval and MBPP, we additionally employ Python post-processing scripts adapted from the dParallel repository chen2025dparallel to handle code execution and pass@1 computation.
Timing and metrics.
We report per-sample averages. End-to-end latency is computed by summing the wall-clock time of the generation routine across all ranks and dividing by the total number of evaluated samples. Total steps denotes the average number of refinement steps actually executed per sample (summed across ranks, then divided by the sample count). Generation length is the average number of valid tokens produced per sample, excluding <endoftext> and any content following a task-specific stop_sequence.
Prompt formatting.
[…]
Appendix B Limitations and Future Directions
Our current training relies on offline, static trajectories. Although we carefully select datasets to limit domain bias, the student can still overfit to the teacher’s priors because supervision does not adapt during training. A straightforward direction is to broaden the data mixture and scale beyond our current $\sim$15k samples, covering more domains.
A complementary path is to move from offline supervision to on-the-fly generation. If the student generates trajectories during training and the teacher verifies or refines them online, we can incorporate on-policy learning to reduce the training-inference discrepancy. In practice, this is currently constrained by DLM sampling speed: slow generation limits the feasibility of closed-loop training without prior trajectory materialization. Improving DLM inference throughput is therefore a prerequisite for large-scale online training.
CDLM’s performance is ultimately bounded by the teacher: a bidirectional DLM distilled into a block-causal student cannot exceed the teacher’s knowledge. To lift this ceiling, distillation from stronger autoregressive (AR) teachers is a natural next step. In this work we intentionally used high-quality, third-party-curated Qwen2.5-7B generations lansechen_easy_2025; lansechen_hard_2025 as ground-truth text, hoping to inject AR knowledge into the student. A more direct, and likely more effective, approach is to distill from an AR teacher: AR models are already scaled up, typically stronger, and much faster at generation than naive DLMs, making large-scale dataset construction easier. Moreover, distillation is well suited to transferring knowledge from a larger, more powerful model to smaller students (here, Dream-7B and LLaDA-8B).